# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v5

@@ 5 Regret Analysis
We begin by analyzing, for each seller $i\in\mathcal{N}$, the convergence of $\widetilde{\bm{\theta}}_{i}$ and $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$ from the exploration phase. An informal overview is given in Section˜5.1, while formal results appear in Appendix˜DSection˜D.1 for $\widetilde{\bm{\theta}}_{i}$ and Section˜D.2 for $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. We then study NE convergence and the regret bound in Section˜5.2.
@@ 5.1 Estimation of the model parameters
In the single-index setting, the linear estimator in equation 10 attains a $\sqrt{n}$ rate when covariates (here, joint prices $\mathbf{p}^{(t)}$) are elliptically distributed (balabdaoui2019least; brillinger2012generalized). We impose this elliptical law during exploration to ensure consistency and, unlike prior work, we derive a concentration inequality for equation 10 (Proposition˜D.3), which underpins our regret bound.
@@
˜5.1 covers Gaussian and, more generally, elliptically symmetric laws with $g(x)=a^{-\gamma x}$ ($a>0$, $\gamma>0$). This yields sub-Gaussian samples, crucial for the concentration result in Proposition˜D.3. We defer to Remark˜D.1 and Remark˜D.2 for a detailed motivation on ˜5.1.
Proposition 5.2 (Informal version of Proposition˜D.3).
@@
Theorem 5.3 (Informal version of Theorem˜D.6).
@@
Suppose that Assumptions 3.1, 3.2, 3.4, 3.6 and 5.1 hold. Then, Algorithm˜1 produces a policy such that, for each $i\in\mathcal{N}$:
@@
where $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in equation 17 and equation 21, respectively.
@@
By Theorem˜5.4, the value of $\xi$ that minimizes the expected regrets $\operatorname{Reg}_{i}(T)$ is the value that equalizes the exponents of $T$ in the regret upper bound: $T^{\xi}+T^{1-\nicefrac{{2\xi}}{{5}}}$ (ignoring logarithmic factors). This yields $\xi=5/7$. Consequently, with this choice of $\xi$, we have that for each $i\in\mathcal{N}$
@@
For $N=1$ (i.e. in a monopolistic setting), our regret upper bound matches that of fan2024policy, who estimate the link function via a kernel-based method. Their result is stated for binary outcomes $y_{i}$, whereas our framework accommodates continuous $y_{i}$; nevertheless, our estimation procedure can be straightforwardly adapted to the binary-response setting, yielding the same regret rate as in the continuous case. Moreover, relative to the kernel approach in fan2024policy, our method offers the advantage that is fully data-driven and requires no bandwidth selection.
@@
As established in Remark˜G.3, if the sellers choose parameters $s_{i}^{\prime}\leq s_{i}$ satisfying $\sup_{i\in\mathcal{N}}\nicefrac{{|1-1/(s_{i}^{\prime}+1)|\,\|\gamma_{i}\|_{1}}}{{\beta_{i}}}<1$, then the convergence rates of the regret and the NE are preserved. This phenomenon is also confirmed empirically in Section˜6.
@@ 6 Numerical Experiments
We evaluate the performance of Algorithm˜1 in markets with $N\in\{2,4,6\}$ sellers. A detailed description of the simulation setup is provided in Appendix˜J; here we present the main setup together with a discussion of the results. For each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$, the price vector $\mathbf{p}^{(t)}$ is supported on $\mathcal{P}=[0,3]^{\times N}$. For every $N$, we examine the convergence behavior as the contraction constant $L_{\bm{\Gamma}}$ varies – specifically, when $L_{\bm{\Gamma}}\approx 0$, $L_{\bm{\Gamma}}\approx 0.5$, and $L_{\bm{\Gamma}}\approx 1$. The link functions are 0-concave and their price sensitivities are generated such that $\bm{\theta}_{i}=(-\beta_{i},\bm{\gamma}_{i})$ has unitary $L^{2}$-norm and satisfies ˜3.6. A fixed point exists and is unique by Lemma˜3.7 and is recovered using a simple root-finding algorithm. The demand noise of each firm follows a uniform distribution on [0.05,0.05][-0.05,0.05]. For every $T\in[100,400,800,1600,3200]$, we apply Algorithm˜1. We independently repeat the simulation $30$ times to obtain average performances and $95\%$ of confidence intervals. In Figure˜6 and Figure˜6 we summarize the results. The panels show: (i) the convergence of the estimators $\bm{\theta}_{i}$ and $\psi_{i}$; (ii) convergence of prices to the Nash equilibrium $\mathbf{p}^{\star}$ and the expected cumulative regret plotted on log–log axes (with fitted slopes) to reveal the empirical rate. Across runs, the observed rates are consistently faster – i.e., exhibit smaller slopes than the theoretical upper bound. Regarding the minimization problem in equation 11 under the class of monotone increasing and $s$-concave functions, we note that designing efficient algorithms specifically tailored to $s$-concavity remains an open and largely unexplored area. Nonetheless, in this work, we leverage a fast optimization solver to approximate the solution. Additionally, in Section˜J.2 we provide simulations under different exploration phases across the sellers and in Section˜J.3 simulations showing how mispecifying $s_{i}$ affects the regret and NE convergence rates.
@@
There are several promising directions for future research. One is to design an algorithm allowing each seller to independently set their exploration length, which is not permitted in our setting as outlined in Remark˜4.1. For this purpose, a key theoretical goal is to obtain uniform convergence for the joint estimator $(\bm{\theta}_{i},\psi_{i})$. While this would not change regrets or NE rates (it will only improve constants), it could yield practical gains: currently, $\bm{\theta}_{i}$ uses $\kappa_{i}\tau$ samples and $\psi_{i}$ the remaining $(1-\kappa_{i})\tau$, whereas a joint estimator would leverage the full $\tau$. Existing results are limited: balabdaoui2019least shows only $L^{2}$ convergence under monotonicity (we additionally require $s_{i}$-concavity), and kuchibhotla2021semiparametric provides uniform convergence under convexity/concavity ($s_{i}=1$) with only $O_{P}$ rates. Our setting instead calls for non-asymptotic supremum-norm concentration for the joint estimator, i.e., tail bounds as in equation 52.
@@
Finally, an ambitious direction is to design algorithms that bypass the classic exploration-exploitation trade-off. Here, each seller $i$ would use the entire history of public prices and private demands $\{(\mathbf{p}^{(s)},y_{i}^{(s)}\}_{s\leq t}$, ensuring $\mathbf{p}^{(T)}\to$ NE as $T\to\infty$. The main challenge is avoiding incomplete learning (keskin2018incomplete), which can yield poor parameter estimates and higher regrets.
@@
In the Supplementary Materials, we provide a .zip archive with all the code required to reproduce the results and figures in this paper. The archive contains three top–level folders whose names match the corresponding figure labels in the PDF. Inside each folder, we include Python notebooks with step-by-step instructions that describe how to run the experiments and regenerate the figure. Running the notebooks as instructed will recreate the plots and save them to the corresponding folders’ directory named “PLOTS".
@@ Appendix A Summary of individual information
In this section, we summarize the information each seller must know to implement Algorithm˜1.
@@
Appendix B Proof of Proposition˜3.5
@@
We have $\varphi_{i}^{\prime}=1+\frac{\psi_{i}^{\prime}\cdot\psi_{i}^{\prime}-\psi_{i}\cdot\psi_{i}^{\prime\prime}}{(\psi_{i}^{\prime})^{2}}=2-\frac{\psi_{i}\cdot\psi_{i}^{\prime\prime}}{(\psi_{i}^{\prime})^{2}}\geq c_{i}$ iff $\psi_{i}\cdot\psi_{i}^{\prime\prime}+(s_{i}-1)(\psi_{i}^{\prime})^{2}\leq 0$ with $s_{i}=c_{i}-1$. The statement follows by Lemma˜B.1.
@@
Proof of Lemma˜B.1.
As for $s=0$ is a known result, we prove it for $s\neq 0$. A function $\psi$ is $s$-concave if and only if $d_{s}\circ\psi$ is concave, where $d_{s}$ is defined in equation 4. Then $\psi$ is $s$-concave if and only if
@@
Appendix C Proof of Lemma˜3.7
@@
First note that, by strong concavity of $p_{i}\to\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})$ (see ˜3.2), the best response map is the projection onto $\mathcal{P}_{i}$ or the value $p_{i}^{*}$ that solves the first order condition. More specifically
@@
where $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $p_{i}^{*}$ solves the first order condition of equation 8, that is
@@
where we that $\mathrm{rev}_{i}(p_{i}|\mathbf{p}_{-i})=p_{i}\psi_{i}(\bm{\theta}_{i}(\mathbf{p}))$. Solving Equation˜15 we have
@@
which is strictly less than $1$ by ˜3.6.
Appendix D Formal estimation of the model parameters in Section˜5.1
@@
In a single-index model, the linear estimator in equation 10 converges to the true $\bm{\theta}_{i}$ at a $\sqrt{n}$ rate provided that the covariates (in this case, the joint prices $\mathbf{p}^{(t)}$) follow an elliptical distribution (balabdaoui2019least; brillinger2012generalized). We adopt the same elliptical assumption on the (joint) distribution of the prices during exploration, to ensure the consistency of the estimator in equation 10. In contrast to earlier work, we establish a concentration inequality for this estimator (Proposition˜D.3), which serves as a key tool in deriving our regret upper bound.
More specifically, in ˜5.1 we assume that $\mathscr{D}=\mathscr{E}(\mathbf{m},\Lambda,g)$, where $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}<c_{\max}<\infty$ and the density generator $g$ is such that $g(x+y)=g(x)g(y)$ for all $x,y\geq 0$, i.e.
@@
˜5.1 holds for a wide range of distributions, including Gaussian distributions and, more generally, all elliptically symmetric distributions for which $g(x)=a^{-\gamma x}$, for $a>0$ and $\gamma>0$. Before describing ˜5.1, we emphasize that selecting an appropriate distribution for the design points (here, the exploration prices) is standard in exploration–exploitation algorithms. This choice is crucial because it ensures that the learner can perform exploration design and obtain a consistent estimator. For example, both fan2024policy and luo2024distribution employ uniformly randomized prices precisely to guarantee consistency of their estimators. We now describe the components of ˜5.1:
@@
    The requirement that exploration prices follow an elliptically symmetric distribution is used to guarantee the consistency of the estimator of $\bm{\theta}_{i}$ in the single-index model. In particular, this assumption is invoked in Lemma˜E.2 to derive the normal equations (the same assumption can be found in balabdaoui2019least and brillinger2012generalized). If the exploration distribution were not elliptically symmetric, the resulting estimator could fail to be consistent.
@@
    As we will discuss in Section˜D.2, the property that $g(x+y)=g(x)g(y)$ (satisfied, for example, by Gaussian distributions) supports the estimation of $\psi_{i}$. It ensures that $\psi_{i,\bm{\theta}}(u)$ (defined in equation 18) depends on $u$ solely through the argument of $\psi_{i}$ rather than that of $g$ (see Equation˜29). This decoupling guarantees that $\psi_{i,\bm{\theta}}$ inherits key properties from $\psi_{i}$, such as smoothness and adherence to the prescribed shape constraints, as highlighted in Proposition˜D.4. In turn, these properties are crucial for establishing consistency of the estimator of $\psi_{i}$.
@@
which is proportional to $\Lambda$. Then, there exists $0<\varsigma_{\min}<\infty$ such that $\varsigma_{\min}\preccurlyeq\Sigma$. These constants will appear in Proposition˜D.3, which proof is provided in Appendix˜E.
@@
Under ˜3.1 and ˜5.1, there exist a positive constant $c$ (that depend solely on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}=\mathbf{p}^{(t)}-\bar{\mathbf{p}}$) such that, for $n_{i}^{(1)}$ sufficiently large, with probability at least $1-2e^{-c\varsigma_{\min}^{2}n_{i}^{(1)}/16}-\frac{2}{n_{i}^{(1)}}$, the estimator in equation 10 satisfies
@@
We now construct an estimator of $\psi_{i,\bm{\theta}}$ and establish, in Theorem˜D.6, a uniform convergence result over any bounded subset $K\subset\mathcal{U}$ (where $\mathcal{U}$ is defined in equation 5). Our approach draws on techniques from mosching2020monotone; dumbgen2004consistency, which study uniform convergence in the context of shape-constrained estimation. A key assumption in these works – one that does not require smoothness of the unknown nonparametric function – is that the density of the design points, $f_{w}$, is bounded away from zero on $\mathcal{U}$. This ensures that the design points $w^{(t)}_{i}\in\mathcal{U}$ are asymptotically dense within every subinterval of $\mathcal{U}$, which in turn is sufficient for establishing uniform convergence of the estimator on such intervals. In our setting, $f_{w}$ is supported on all of $\mathbb{R}$, and thus it is bounded away from zero on any bounded set $K$. Consequently, the uniform convergence result applies to any bounded subset $K\subset\mathcal{U}$.
Before establishing the convergence result, we must first characterize the shape constraints satisfied by the family of functions $\{\psi_{i,\bm{\theta}}(\cdot):\bm{\theta}\in\mathbb{S}_{N-1}\}$. In Proposition˜D.4 (with proof deferred to Appendix˜F), we derive a closed-form expression for $\psi_{i,\bm{\theta}}(\cdot)$ and show that both the monotonicity and the $s_{i}=c_{i}-1$-concavity of the original function $\psi_{i}$ are preserved by $\psi_{i,\bm{\theta}}$, uniformly over $\bm{\theta}\in\mathbb{S}_{N-1}$.
@@
Under ˜3.1 and ˜5.1 we have the following properties:
@@
    If also ˜3.4 holds (i.e. $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$ as it arises from Proposition˜3.5) then $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for any $\bm{\theta}\in\mathbb{S}_{N-1}$.
By definition of $s$-concavity, from point (5) in Proposition˜D.4 we have that for every $\bm{\theta}\in\mathbb{S}_{N-1}$ there exists a concave function $\phi_{i,\bm{\theta}}$ that $\psi_{i,\bm{\theta}}=h_{s_{i}}\circ\phi_{i,\bm{\theta}}$, where $h_{s_{i}}$ is defined in equation 4, and $s_{i}=c_{i}-1$ is known by ˜3.4. We could then estimate $\phi_{i,\bm{\theta}}(u)$ using LSE under concavity restriction, that is
@@
where we recall $w^{(t)}_{i}\triangleq\langle\bm{\theta},\mathbf{p}^{(t)}\rangle$. The interval $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is such that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$. Note that $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is known because $\underline{B}_{\psi_{i}}$ and $\bar{B}_{\psi_{i}}$ are known by ˜3.1 (recll that, by point (2) in Proposition˜D.4, the bounds $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are inherited from $\psi_{i}$ to $\psi_{i,\bm{\theta}}$). We want to highlight that for any fixed $\bm{\theta}\in\mathbb{S}_{N-1}$, the problem (19) is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi_{i}(w_{1}^{(t)}),\phi_{i}(w_{2}^{(t)}),\dots,\phi_{i}(w_{n}^{(t)}))\in[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$. A full discussion on the optimization problem (19) can be found in Appendix˜H.

### v6

@@ 5 Regret Analysis
We begin by analyzing, for each seller $i\in\mathcal{N}$, the convergence of $\widetilde{\bm{\theta}}_{i}$ and $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$ from the exploration phase. An informal overview is given in Section 5.1, while formal results appear in Appendix DSection D.1 for $\widetilde{\bm{\theta}}_{i}$ and Section D.2 for $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. We then study NE convergence and the regret bound in Section 5.2.
@@ 5.1 Estimation of the model parameters
In the single-index setting, the linear estimator in (10) attains a $\sqrt{n}$ rate when covariates (here, joint prices $\mathbf{p}^{(t)}$) are elliptically distributed (Balabdaoui et al., 2019; Brillinger, 2012). We impose this elliptical law during exploration to ensure consistency and, unlike prior work, we derive a concentration inequality for (10) (Proposition D.3), which underpins our regret bound.
@@
5.1 covers Gaussian and, more generally, elliptically symmetric laws with $g(x)=a^{-\gamma x}$ ($a>0$, $\gamma>0$). This yields sub-Gaussian samples, crucial for the concentration result in Proposition D.3. We defer to Remark D.1 and Remark D.2 for a detailed motivation on 5.1.
Proposition 5.2 (Informal version of Proposition D.3).
@@
Theorem 5.3 (Informal version of Theorem D.6).
@@
Suppose that Assumptions 3.1, 3.2, 3.4, 3.6 and 5.1 hold. Then, Algorithm 1 produces a policy such that, for each $i\in\mathcal{N}$:
@@
where $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in (17) and (21), respectively.
@@
By Theorem 5.4, the value of $\xi$ that minimizes the expected regrets $\operatorname{Reg}_{i}(T)$ is the value that equalizes the exponents of $T$ in the regret upper bound: $T^{\xi}+T^{1-\nicefrac{{2\xi}}{{5}}}$ (ignoring logarithmic factors). This yields $\xi=5/7$. Consequently, with this choice of $\xi$, we have that for each $i\in\mathcal{N}$
@@
For $N=1$ (i.e. in a monopolistic setting), our regret upper bound matches that of Fan et al. (2024), who estimate the link function via a kernel-based method. Their result is stated for binary outcomes $y_{i}$, whereas our framework accommodates continuous $y_{i}$; nevertheless, our estimation procedure can be straightforwardly adapted to the binary-response setting, yielding the same regret rate as in the continuous case. Moreover, relative to the kernel approach in Fan et al. (2024), our method offers the advantage that is fully data-driven and requires no bandwidth selection.
@@
As established in Remark G.3, if the sellers choose parameters $s_{i}^{\prime}\leq s_{i}$ satisfying $\sup_{i\in\mathcal{N}}\nicefrac{{|1-1/(s_{i}^{\prime}+1)|\,\|\gamma_{i}\|_{1}}}{{\beta_{i}}}<1$, then the convergence rates of the regret and the NE are preserved. This phenomenon is also confirmed empirically in Section 6.
@@ 6 Numerical Experiments
We evaluate the performance of Algorithm 1 in markets with $N\in\{2,4,6\}$ sellers. A detailed description of the simulation setup is provided in Appendix J; here we present the main setup together with a discussion of the results. For each seller $i\in\mathcal{N}=\{1,2,\dots,N\}$, the price vector $\mathbf{p}^{(t)}$ is supported on $\mathcal{P}=[0,3]^{\times N}$. For every $N$, we examine the convergence behavior as the contraction constant $L_{\bm{\Gamma}}$ varies – specifically, when $L_{\bm{\Gamma}}\approx 0$, $L_{\bm{\Gamma}}\approx 0.5$, and $L_{\bm{\Gamma}}\approx 1$. The link functions are 0-concave and their price sensitivities are generated such that $\bm{\theta}_{i}=(-\beta_{i},\bm{\gamma}_{i})$ has unitary $L^{2}$-norm and satisfies 3.6. A fixed point exists and is unique by Lemma 3.7 and is recovered using a simple root-finding algorithm. The demand noise of each firm follows a uniform distribution on [0.05,0.05][-0.05,0.05]. For every $T\in[100,400,800,1600,3200]$, we apply Algorithm 1. We independently repeat the simulation $30$ times to obtain average performances and $95\%$ of confidence intervals. In Figure 6 and Figure 6 we summarize the results. The panels show: (i) the convergence of the estimators $\bm{\theta}_{i}$ and $\psi_{i}$; (ii) convergence of prices to the Nash equilibrium $\mathbf{p}^{\star}$ and the expected cumulative regret plotted on log–log axes (with fitted slopes) to reveal the empirical rate. Across runs, the observed rates are consistently faster – i.e., exhibit smaller slopes than the theoretical upper bound. Regarding the minimization problem in (11) under the class of monotone increasing and $s$-concave functions, we note that designing efficient algorithms specifically tailored to $s$-concavity remains an open and largely unexplored area. Nonetheless, in this work, we leverage a fast optimization solver to approximate the solution. Additionally, in Section J.2 we provide simulations under different exploration phases across the sellers and in Section J.3 simulations showing how mispecifying $s_{i}$ affects the regret and NE convergence rates.
@@
There are several promising directions for future research. One is to design an algorithm allowing each seller to independently set their exploration length, which is not permitted in our setting as outlined in Remark 4.1. For this purpose, a key theoretical goal is to obtain uniform convergence for the joint estimator $(\bm{\theta}_{i},\psi_{i})$. While this would not change regrets or NE rates (it will only improve constants), it could yield practical gains: currently, $\bm{\theta}_{i}$ uses $\kappa_{i}\tau$ samples and $\psi_{i}$ the remaining $(1-\kappa_{i})\tau$, whereas a joint estimator would leverage the full $\tau$. Existing results are limited: Balabdaoui et al. (2019) shows only $L^{2}$ convergence under monotonicity (we additionally require $s_{i}$-concavity), and Kuchibhotla et al. (2021) provides uniform convergence under convexity/concavity ($s_{i}=1$) with only $O_{P}$ rates. Our setting instead calls for non-asymptotic supremum-norm concentration for the joint estimator, i.e., tail bounds as in (52).
@@
Finally, an ambitious direction is to design algorithms that bypass the classic exploration-exploitation trade-off. Here, each seller $i$ would use the entire history of public prices and private demands $\{(\mathbf{p}^{(s)},y_{i}^{(s)}\}_{s\leq t}$, ensuring $\mathbf{p}^{(T)}\to$ NE as $T\to\infty$. The main challenge is avoiding incomplete learning (Keskin and Zeevi, 2018), which can yield poor parameter estimates and higher regrets.
@@
In the Supplementary Materials, we provide a .zip archive with all the code required to reproduce the results and figures in this paper. The archive contains three top–level folders whose names match the corresponding figure labels in the PDF. Inside each folder, we include Python notebooks with step-by-step instructions that describe how to run the experiments and regenerate the figure. Running the notebooks as instructed will recreate the plots and save them to the corresponding folders’ directory named “PLOTS”.
@@ Appendix A Summary of individual information
In this section, we summarize the information each seller must know to implement Algorithm 1.
@@
Appendix B Proof of Proposition 3.5
@@
We have $\varphi_{i}^{\prime}=1+\frac{\psi_{i}^{\prime}\cdot\psi_{i}^{\prime}-\psi_{i}\cdot\psi_{i}^{\prime\prime}}{(\psi_{i}^{\prime})^{2}}=2-\frac{\psi_{i}\cdot\psi_{i}^{\prime\prime}}{(\psi_{i}^{\prime})^{2}}\geq c_{i}$ iff $\psi_{i}\cdot\psi_{i}^{\prime\prime}+(s_{i}-1)(\psi_{i}^{\prime})^{2}\leq 0$ with $s_{i}=c_{i}-1$. The statement follows by Lemma B.1.
@@
Proof of Lemma B.1.
As for $s=0$ is a known result, we prove it for $s\neq 0$. A function $\psi$ is $s$-concave if and only if $d_{s}\circ\psi$ is concave, where $d_{s}$ is defined in (4). Then $\psi$ is $s$-concave if and only if
@@
Appendix C Proof of Lemma 3.7
@@
First note that, by strong concavity of $p_{i}\to\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})$ (see 3.2), the best response map is the projection onto $\mathcal{P}_{i}$ or the value $p_{i}^{*}$ that solves the first order condition. More specifically
@@
where $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $p_{i}^{*}$ solves the first order condition of (8), that is
@@
where we that $\mathrm{rev}_{i}(p_{i}|\mathbf{p}_{-i})=p_{i}\psi_{i}(\bm{\theta}_{i}(\mathbf{p}))$. Solving Equation 15 we have
@@
which is strictly less than $1$ by 3.6.
Appendix D Formal estimation of the model parameters in Section 5.1
@@
In a single-index model, the linear estimator in (10) converges to the true $\bm{\theta}_{i}$ at a $\sqrt{n}$ rate provided that the covariates (in this case, the joint prices $\mathbf{p}^{(t)}$) follow an elliptical distribution (Balabdaoui et al., 2019; Brillinger, 2012). We adopt the same elliptical assumption on the (joint) distribution of the prices during exploration, to ensure the consistency of the estimator in (10). In contrast to earlier work, we establish a concentration inequality for this estimator (Proposition D.3), which serves as a key tool in deriving our regret upper bound.
More specifically, in 5.1 we assume that $\mathscr{D}=\mathscr{E}(\mathbf{m},\Lambda,g)$, where $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}<c_{\max}<\infty$ and the density generator $g$ is such that $g(x+y)=g(x)g(y)$ for all $x,y\geq 0$, i.e.
@@
5.1 holds for a wide range of distributions, including Gaussian distributions and, more generally, all elliptically symmetric distributions for which $g(x)=a^{-\gamma x}$, for $a>0$ and $\gamma>0$. Before describing 5.1, we emphasize that selecting an appropriate distribution for the design points (here, the exploration prices) is standard in exploration–exploitation algorithms. This choice is crucial because it ensures that the learner can perform exploration design and obtain a consistent estimator. For example, both Fan et al. (2024) and Luo et al. (2024) employ uniformly randomized prices precisely to guarantee consistency of their estimators. We now describe the components of 5.1:
@@
    The requirement that exploration prices follow an elliptically symmetric distribution is used to guarantee the consistency of the estimator of $\bm{\theta}_{i}$ in the single-index model. In particular, this assumption is invoked in Lemma E.2 to derive the normal equations (the same assumption can be found in Balabdaoui et al. (2019) and Brillinger (2012)). If the exploration distribution were not elliptically symmetric, the resulting estimator could fail to be consistent.
@@
    As we will discuss in Section D.2, the property that $g(x+y)=g(x)g(y)$ (satisfied, for example, by Gaussian distributions) supports the estimation of $\psi_{i}$. It ensures that $\psi_{i,\bm{\theta}}(u)$ (defined in (18)) depends on $u$ solely through the argument of $\psi_{i}$ rather than that of $g$ (see Equation 29). This decoupling guarantees that $\psi_{i,\bm{\theta}}$ inherits key properties from $\psi_{i}$, such as smoothness and adherence to the prescribed shape constraints, as highlighted in Proposition D.4. In turn, these properties are crucial for establishing consistency of the estimator of $\psi_{i}$.
@@
which is proportional to $\Lambda$. Then, there exists $0<\varsigma_{\min}<\infty$ such that $\varsigma_{\min}\preccurlyeq\Sigma$. These constants will appear in Proposition D.3, which proof is provided in Appendix E.
@@
Under 3.1 and 5.1, there exist a positive constant $c$ (that depend solely on the variance proxy $\sigma_{x}$ of the sub-Gaussian random variable $\mathbf{x}^{(t)}=\mathbf{p}^{(t)}-\bar{\mathbf{p}}$) such that, for $n_{i}^{(1)}$ sufficiently large, with probability at least $1-2e^{-c\varsigma_{\min}^{2}n_{i}^{(1)}/16}-\frac{2}{n_{i}^{(1)}}$, the estimator in (10) satisfies
@@
We now construct an estimator of $\psi_{i,\bm{\theta}}$ and establish, in Theorem D.6, a uniform convergence result over any bounded subset $K\subset\mathcal{U}$ (where $\mathcal{U}$ is defined in (5)). Our approach draws on techniques from Mösching and Dümbgen (2020); Dümbgen et al. (2004), which study uniform convergence in the context of shape-constrained estimation. A key assumption in these works – one that does not require smoothness of the unknown nonparametric function – is that the density of the design points, $f_{w}$, is bounded away from zero on $\mathcal{U}$. This ensures that the design points $w^{(t)}_{i}\in\mathcal{U}$ are asymptotically dense within every subinterval of $\mathcal{U}$, which in turn is sufficient for establishing uniform convergence of the estimator on such intervals. In our setting, $f_{w}$ is supported on all of $\mathbb{R}$, and thus it is bounded away from zero on any bounded set $K$. Consequently, the uniform convergence result applies to any bounded subset $K\subset\mathcal{U}$.
Before establishing the convergence result, we must first characterize the shape constraints satisfied by the family of functions $\{\psi_{i,\bm{\theta}}(\cdot):\bm{\theta}\in\mathbb{S}_{N-1}\}$. In Proposition D.4 (with proof deferred to Appendix F), we derive a closed-form expression for $\psi_{i,\bm{\theta}}(\cdot)$ and show that both the monotonicity and the $s_{i}=c_{i}-1$-concavity of the original function $\psi_{i}$ are preserved by $\psi_{i,\bm{\theta}}$, uniformly over $\bm{\theta}\in\mathbb{S}_{N-1}$.
@@
Under 3.1 and 5.1 we have the following properties:
@@
    If also 3.4 holds (i.e. $\psi_{i}$ is $s_{i}$-concave for some $s_{i}>-1$ as it arises from Proposition 3.5) then $\psi_{i,\bm{\theta}}$ is $s_{i}$-concave for any $\bm{\theta}\in\mathbb{S}_{N-1}$.
By definition of $s$-concavity, from point (5) in Proposition D.4 we have that for every $\bm{\theta}\in\mathbb{S}_{N-1}$ there exists a concave function $\phi_{i,\bm{\theta}}$ that $\psi_{i,\bm{\theta}}=h_{s_{i}}\circ\phi_{i,\bm{\theta}}$, where $h_{s_{i}}$ is defined in (4), and $s_{i}=c_{i}-1$ is known by 3.4. We could then estimate $\phi_{i,\bm{\theta}}(u)$ using LSE under concavity restriction, that is
@@
where we recall $w^{(t)}_{i}\triangleq\langle\bm{\theta},\mathbf{p}^{(t)}\rangle$. The interval $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is such that $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}]=h_{s_{i}}([\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}])$. Note that $\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$ is known because $\underline{B}_{\psi_{i}}$ and $\bar{B}_{\psi_{i}}$ are known by 3.1 (recll that, by point (2) in Proposition D.4, the bounds $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are inherited from $\psi_{i}$ to $\psi_{i,\bm{\theta}}$). We want to highlight that for any fixed $\bm{\theta}\in\mathbb{S}_{N-1}$, the problem (19) is not infinite dimensional, indeed we only need to recover $n$ variables $(\phi_{i}(w_{1}^{(t)}),\phi_{i}(w_{2}^{(t)}),\dots,\phi_{i}(w_{n}^{(t)}))\in[\underline{B}_{\phi_{i}},\bar{B}_{\phi_{i}}]$. A full discussion on the optimization problem (19) can be found in Appendix H.